# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ IV Scalability of synchronized multiphoton generation
We are now in position to discuss the most important results of the present work: the scalability performance of the multiphoton emitter. To do so, we analyze the $N$-photon generation probability $P_{N}$ and the demultiplexing error $D_{N}$ as a function of $N$ and the various system parameters. From these results we demonstrate the feasibility to synchronize $N\sim 100-1000$ emitters with high $N$-photon generation rates $C_{N}=RP_{N}$.
@@
We calculate $P_{N}$ in Eq. (20) by solving the master equation (12), which includes all qubit-qubit correlations induced by the cavity ($\sim J_{\rm cm},\gamma_{\rm cm}$), as well as the decoherence effects ($\sim\gamma_{\phi},\gamma_{\rm loss}$). The Hilbert space of the system grows exponentially as $2^{N}$, but using quantum trajectories (QT) [cf. Appendix B.2] we can estimate the photon statistics of the multiphoton emission and thereby $P_{N}$ up to moderately large numbers of SPSs. Figure. 6(a) displays $P_{N}$ as a function of $N$ computed from an average over $M=6000$ trajectories and considering the four parameter sets of Table 1. In the four cases and up to $N=12$, we numerically confirm that the SPSs are nearly perfectly synchronized and independent, satisfying $P_{N}\approx({\cal P}_{1})^{N}$ within the the statistical error $\Delta P_{N}\sim 10^{-3}$ [cf. black lines in Fig. (6)]. The reduction of $N$-photon generation probability with $N$ is thus mainly limited by the imperfections in the single-photon efficiency ${\cal P}_{1}$ and not by the synchronizing and demultiplexing scheme. For the decoherence-free parameters A, we predict a $12$-photon efficiency as high as $P_{12}\approx 0.97$, and when including realistic decoherence sources as in parameter set C [cf. inset in Fig. 6(a)], this reduces only to $P_{12}\approx 0.88$. For a pulsed excitation of the SPSs with a repetition rate $R=\gamma/15$ (as in Secs. III.4-III.5), these high efficiencies imply $12$-photon generation rates $C_{12}=RP_{12}$ of $\sim 0.1-1$MHz [cf. Fig. 6(b) and inset]. Since parameter set D has the largest antenna decay $\gamma/2\pi\sim 3$MHz, it shows the largest $N$-photon generation rates $C_{N}$ despite having the lowest efficiencies $P_{N}$ and the largest decoherence rates.
@@
To quantify more precisely the deviations from the ideal scaling $P_{N}\approx({\cal P}_{1})^{N}$ in Figs. 6(a)-(b), we compute the $N$-photon dependence or demultiplexing error $D_{N}$ as defined in Eq. (21). For $N=2$ we know this error is as small as $D_{2}\sim 10^{-7}$ [cf. Figs. 3] and therefore the QT calculations with an uncertainty $\Delta P_{N}\sim 10^{-3}$ do not provide enough precision. To address this problem, we developed a photon counting approach based on the master equation [cf. Appendix B.1], which does not suffer from any statistical uncertainty. In this alternative method we simulate the photon counters at each antenna by an additional two-level system. This increases the Hilbert space dimension to $2^{2N}$ and thus limits the numerical computations to a maximum of $N\sim 5$ emitters.
In the inset of Fig. 6(c) we show the results of $D_{N}$ as a function $N$ for the four parameter sets of Table 1, and up to $N=5$. We confirm that $D_{N}<10^{-5}$ is well below the QT precision, and most importantly, we observe that the data is very well approximated by the quadratic fit
$D_{N}=\epsilon N(N-1),$ (27)
where the scaling factor $\epsilon\ll 1$ depends on the system parameters [cf. inset of Fig. 6(c)]. To analyze how the cavity-induced correlations influence $\epsilon$ and thus the scalability of the multiphoton scheme, we vary the system parameters around parameter set B and extract $\epsilon$ from fits to numerical calculations of $D_{N}$ up to $N=5$. The results of this analysis are represented by the data points in Fig. 7(a), which correspond to independent sweeps over each of the system parameters. Remarkably, the apparently complicated behaviour of the system is very well approximated by the analytical formula [cf. solid lines in Fig. 7(a)],
$\epsilon\approx(1.42)\frac{g^{4}}{\gamma^{2}}\frac{\kappa}{\left[(\kappa/2)^{2}+\Delta^{2}\right]^{3/2}},$ (28)
which manifests the dominant dependence on $g$, $\gamma$, $\kappa$, and $\Delta$, and omits the marginal effects of the effective driving $\Omega_{\rm cm}$, and the decoherence rates $\gamma_{\rm loss}$ and $\gamma_{\phi}$. This confirms our intuition that the synchronization and demultiplexing performance of our scheme is limited exclusively by residual cavity-induced correlations, provided $J_{\rm cm},\gamma_{\rm cm}\ll\gamma$ [cf. Eq. (18)].
Although Eq. (28) was extracted from a variation of parameters around set B of Table 1, this expression accurately predicts the scaling factors $\epsilon$ obtained from the fits on all parameters sets A-D in Fig. 6(c) with an error smaller than $8\%$. Interestingly, we can exploit Eq. (28) to predict new parameter sets whose demultiplexing error grows quickly to $D_{N}\gtrsim 1$ for moderate system sizes $N\lesssim 10$. Setups with these characteristics can then be used to study the validity limits of the approximated $D_{N}$ scaling in Eq. (27). In Fig. 7(b) we illustrate this analysis by calculating $D_{N}$ for two parameter sets similar to B, but: (i) with a $10$-fold coupling enhancement $g/2\pi=2$MHz (blue data), and (ii) with a $1/2$ reduction of decay $\gamma/2\pi=0.2$MHz in addition to the same $10$-fold increase in $g$ (red data). We perform $D_{N}=\epsilon N(N-1)$ fits to both data sets up to $N=5$, and we clearly observe the on-set of deviations from this scaling only when $D_{N}\gtrsim 1$ [cf. solid red line in Fig. 7]. For $D_{N}\ll 1$, both data sets show excellent agreement with the quadratic scaling (27), and this constitutes the range of validity of this approximation.
Figure 7: Limits on the scalability of the synchronized multiphoton generation. (a) Scaling factor $\epsilon$ as a function of various system parameters. The dimensionless variable $x=1$ corresponds to the parameters of set B in Table 1, which are denoted by $\tilde{g}$, $\tilde{\gamma}$, $\tilde{\Delta}$, $\tilde{\kappa}$, $\tilde{\Omega}_{\rm cm}$, $\tilde{\gamma}_{\rm loss}$, $\tilde{\gamma}_{\phi}$, and in this plot we change each of them from $25\%$ to $400\%$. We use these results to numerically extract Eq. (28). (b) Demultiplexing error $D_{N}$ calculated up to $N=10$ with QT averaged over $M=10^{4}$ trajectories. Both parameter sets are similar to B, except for $g/2\pi=2$MHz (blue data) and $g/2\pi,\gamma/2\pi]=[2,0.2]$MHz (red data). Solid lines correspond to fits of $D_{N}=\epsilon N(N-1)$ up to $N=5$, which result in the $\epsilon$ factors shown in the legend. For both data sets we observe that the quadratic scaling is a good approximation as long as $D_{N}\ll 1$, confirming the expected limit of validity.
@@ V Inhomogeneity effects
So far, we have analyzed the performance of the scheme assuming a perfectly homogeneous setup. In large scale implementations, however, the SPS parameters will unavoidably present some degree of inhomogeneity. In this last section we analyze the impact of these imperfections.
@@
Here, $\Delta y$ denotes the standard deviation associated to the disorder on each qubit parameter $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$.
To quantify the impact of each of the inhomogeneities $\delta y_{j}$ on the performance of the multiphoton emitter, we compute the average $N$-photon efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ as
@@
where the probabilities $P_{N}[m]_{y}$ are calculated with the master equation method of Appendix B.1 for each realization $m=1,\dots,M$ of the independent disorder $y$. The results of these computations for $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ in the case of disorder on the antenna couplings $\Delta\gamma$, as well as in the decoherence rates $\Delta\gamma_{\phi}$ and $\Delta\gamma_{\rm loss}$, are shown in Fig. (8). We perform calculations up to $N=4$ slightly inhomogeneous photon sources, for parameters B and C, and we do not observe any detrimental effect up to a disorder strength of $10\%$ of the average values [cf. Fig. 8(a)-(b)]. This is not surprising in the case of inhomogeneous decoherence rates $\gamma_{\phi}^{j}$ and $\gamma_{\rm loss}^{j}$ since their effects have nothing to do with the synchronization dynamics and are very marginally small anyways $\gamma_{\phi}^{j},\gamma_{\rm loss}^{j}\ll\gamma$. In the case of inhomogeneous antenna couplings $\gamma_{j}=\gamma+\delta\gamma_{j}$, they imply slightly different emission time scales for the emitted single-photons,
@@
Figure 8: Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{y}$, as a function of disorder in antenna decay $y=\Delta\gamma/\gamma$, qubit dephasing $y=\Delta\gamma_{\phi}/\gamma_{\phi}$, and waveguide loss $y=\Delta\gamma_{\rm loss}/\gamma_{\rm loss}$. (a)-(b) Up to a disorder strength of $10\%$, we do not observe any impact of these three disorder sources on the synchronization and demultiplexing dynamics, as calculated up to $N=4$ for parameter sets B and C, respectively. All results are obtained from averaging $M=10^{3}$ realizations for each type of disorder.
On the other hand, inhomogeneities in the qubit frequencies $\delta\omega^{j}_{q}$ and in the couplings $\delta g_{j}$ are much more harmful for the multiphoton synchronization performance and therefore they need to be controlled more precisely in an experimental implementation of the device. Inhomogeneous couplings $g_{j}$, in particular, induce different cavity-mediated driving strengths $\Omega^{j}_{\rm cm}=\Omega_{\rm cm}+|\alpha_{\rm ss}|\delta g_{j}$, and therefore different times to realize an exact $\pi$-pulse on each qubit $\sim\pi/(2\Omega_{\rm cm}^{j})$. Since we control only the global duration of the cavity pulse, we optimally set it to the average $\pi$-pulse time $T=\pi/(2\Omega_{\rm cm})$, but this unavoidably leads to slightly different probabilities $P_{e}^{j}$ of preparing the excited states $|e\rangle_{j}$ on each qubit. Explicitly, we have
@@
and therefore a low coupling disorder $\delta g_{j}\ll g$ is required for a high performance of the synchronized multiphoton emitter. To precisely quantity the deviation from the ideal photon independence condition, we define the average dependence or demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{y}$ over disorder $y$ as
@@
with ${\cal P}_{1}^{j}[m]_{y}$ calculated from $m=1,\dots,M$ realizations of the disorder $y$ and for each of the $j=1,\dots,N$ SPSs in the setup. In Figs. 9(a)-(b) we compute $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=4$ for parameter set B, and disorder strengths $\Delta g/g=0\%$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green). To achieve reasonable statistical errors $\Delta D_{N}\sim 10^{-6}$, these calculations require up to $M=3\cdot 10^{5}$ realizations. We observe that the $N$-photon efficiency and dependence, $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$, are minimally modified as long as $\delta g_{j}/g\lesssim 1\%$. Fortunately, current superconducting circuit technology allows for less than $1\%$ of geometrical errors in device fabrication, which is thus enough to suppress the disorder effects in the design of capacitive qubit-cavity couplings $g_{j}$.
Figure 9: Impact of inhomogeneities in qubit-cavity couplings $\delta g_{j}$ and qubit frequencies $\delta\omega_{j}$ on the performance of the synchronized multiphoton emitter. (a) Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ up to $N=4$ for coupling disorder strengths $\Delta g/g=0$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green). (b) Average demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=3$, for the same disorder values as in (a). (c)-(d) Analogous calculations as in (a)-(b), but considering $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ with disorder in the qubit frequencies $\delta\omega^{j}_{q}$. We consider disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green). All calculations are performed using the parameter set B of Table 1 and up to $M=3\cdot 10^{5}$ realizations of each type of disorder.

### v3

@@ IV Scalability of synchronized multiphoton generation
Figure 6: Scalability performance of the synchronized multiphoton generation scheme. (a) NN-photon generation efficiency $P_{N}$ as a function of $N$, for parameter sets A (green), B (blue), C (red), and D (grey), as given in Table 1. Data points up to $N=12$ correspond to QT calculations averaged over $M=6000$ trajectories, which show excellent agreement (within the statistical error $\Delta P_{N}\sim 10^{-3}$) with the prediction for perfectly independent and synchronized SPSs, i.e. $P_{N}\approx({\cal P}_{1})^{N}$. [cf. black lines and inset]. (b) NN-photon generation rate $C_{N}=RP_{N}$ for pulsed excitation with repetition rate $R=\gamma/15$. Despite parameter set B shows the lowest efficiency $P_{N}$, the faster repetition rate $R\approx 1.26$MHz allows for appreciably higher $C_{N}$. (c) Demultiplexing error $D_{N}$ up to $N=5$, and quadratic fit $D_{N}=\epsilon N(N-1)$ (cf. solid lines and inset). The fitted scaling factor depends on the system parameters and reads $\epsilon_{A}=1.3\times 10^{-4}$, $\epsilon_{B}=1.2\times 10^{-5}$, $\epsilon_{C}=6.7\times 10^{-7}$, and $\epsilon_{D}=1.0\times 10^{-7}$ for parameter sets A, B, C, and D, respectively. The extrapolation of all these calculations is valid as long as $D_{N}\ll 1$.
@@
We are now in a position to discuss the most important results of the present work: the scalability performance of the multiphoton emitter. To do so, we analyze the $N$-photon generation probability $P_{N}$ and the demultiplexing error $D_{N}$ as a function of $N$, for the four parameter sets of Table 1.
@@
We calculate $P_{N}$ in Eq. (21) by solving the master equation (13), which includes all qubit-qubit correlations induced by the cavity ($\sim J_{\rm cm},\gamma_{\rm cm}$), as well as the decoherence effects ($\sim\gamma_{\phi},\gamma_{\rm loss}$). The Hilbert space of the system grows exponentially as $2^{N}$, but using quantum trajectories (QT) [cf. Appendix B.2] we can estimate the photon statistics of the multiphoton emission and thereby $P_{N}$ up to moderately large numbers of SPSs. Figure. 6(a) displays $P_{N}$ as a function of $N$ computed from an average over $M=6000$ trajectories and considering the four parameter sets of Table 1. In all these cases and up to $N=12$, we numerically confirm that the SPSs are nearly perfectly synchronized and independent, satisfying $P_{N}\approx({\cal P}_{1})^{N}$ within the statistical error $\Delta P_{N}\sim 10^{-3}$ [cf. black lines in Fig. (6)]. The reduction of $N$-photon generation probability with $N$ is thus mainly limited by the imperfections in the single-photon efficiency ${\cal P}_{1}$ and not by the synchronizing and demultiplexing scheme. For the state-of-the-art parameters A and B, we predict a $10$-photon efficiency of $P_{10}\approx 0.90$ and $0.83$, respectively [cf. inset in Fig. 6(a)]. The reduction of efficiency for B is mainly due to the larger $\gamma$, but this also allows it to have a much larger $10$-photon emission rate of $C_{10}=1.0$ MHz compared to $C_{10}=0.4$ MHz for A [cf. Fig. 6(b)]. The long-term parameters C and D present similar $N$-photon emission rates, but much larger efficiencies of $P_{10}\approx 0.94$ and $0.91$, respectively.
@@
To quantify more precisely the deviations from the ideal scaling $P_{N}\approx({\cal P}_{1})^{N}$ in Figs. 6(a)-(b), we compute the $N$-photon dependence or demultiplexing error $D_{N}$ as defined in Eq. (22). For $N=2$ we know this error is as small as $D_{2}\sim{\color[rgb]{0,0,0}10^{-4}}$ [cf. Fig. 3] and therefore the QT calculations with an uncertainty $\Delta P_{N}\sim 10^{-3}$ do not provide enough precision. To address this problem, we developed a photon counting approach based on the master equation [cf. Appendix B.1], which does not suffer from any statistical uncertainty. In this alternative method, we simulate the photon counters at each antenna by an additional two-level system. This increases the Hilbert space dimension to $2^{2N}$ and thus limits the numerical computations to a maximum of $N\sim 5$ emitters. In the inset of Fig. 6(c) we show the results of $D_{N}$ as a function $N$ for the four parameter sets of Table 1, and up to $N=5$. Importantly, we observe that the data is very well approximated by the quadratic fit $D_{N}=\epsilon N(N-1)$ stated in Eq. (4), with a scaling factor $\epsilon\ll 1$ that depends on the system parameters. For parameter sets A, B, C, and D, we obtain $\epsilon_{A}=1.3\times 10^{-4}$, $\epsilon_{B}=1.2\times 10^{-5}$, $\epsilon_{C}=6.7\times 10^{-7}$, and $\epsilon_{D}=1.0\times 10^{-7}$, respectively [cf. inset of Fig. 6(c)]. Notice that $\epsilon$ is smaller when the conditions $\gamma_{\rm cm},J_{\rm cm}\ll\gamma$ are better fulfilled overall, confirming that the synchronization performance is limited mainly by residual cavity-induced correlations [cf. Eq. (19)].
We will now analyze the validity limits of the quadratic scaling $D_{N}=\epsilon N(N-1)$. This is hard to do for the optimized parameter A-D because the deviation is expected for $N\gg 10$ due to the low correlation errors $\epsilon$. However, we can reduce $\gamma$ and/or increase $g$ to observe the growth of the demultiplexing error with $N$ up to $D_{N}\gtrsim 1$ with moderate system sizes $N\lesssim 10$. In Fig. 7, we illustrate this analysis by calculating $D_{N}$ for two modified parameter sets similar to A but with a 10-fold lower antenna decay $\gamma/2\pi=0.1$ MHz (blue diamonds), and with that lower decay and also larger coupling $\gamma/2\pi,g/2\pi]=[0.1,6]$MHz (red crosses). We also include the results for the standard parameter set A as a reference (blue circles). We perform quadratic fits $D_{N}=\epsilon N(N-1)$ to the three data sets up to $N=5$, obtaining increasing correlation errors: $\epsilon=1.4\times 10^{-4}$ (green), $\epsilon=8.5\times 10^{-3}$ (blue), and $\epsilon=3.5\times 10^{-2}$ (red). Remarkably, we indeed observe the on-set of deviations from the quadratic scaling only when $D_{N}\gtrsim 1$, which occurs for the red data at $N\sim 7$ [cf. solid red line in Fig. 7]. For $D_{N}\ll 1$, all these data sets show excellent agreement with the quadratic scaling $D_{N}=\epsilon N(N-1)$, and this constitutes the range of validity of this approximation.
Figure 7: Demultiplexing error $D_{N}$ calculated up to $N=9$ with QT averaged over $M=10^{4}$ trajectories. Green circles correspond to the parameter set A of Table 1, whereas the other two data sets only differ in having $\gamma/2\pi=0.1$MHz (blue diamonds) and $\gamma/2\pi,g/2\pi]=[0.1,6]$MHz (red crosses). Solid lines correspond to fits of $D_{N}=\epsilon N(N-1)$ up to $N=5$, which result in the $\epsilon$ factors shown in the legend. We observe that the quadratic scaling deviates only for $D_{N}\gtrsim 1$, confirming the validity range $D_{N}\ll 1$.
@@ V Inhomogeneity effects
So far, we have analyzed the performance of the scheme assuming a perfectly homogeneous setup. In large-scale implementations, however, the SPS parameters will unavoidably present some degree of inhomogeneity. In this section, we analyze the impact of these imperfections.
@@
Here, $\Delta y$ denotes the standard deviation associated with the disorder on each qubit parameter $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$. To quantify the impact of each of the inhomogeneities $\delta y_{j}$ on the performance of the multiphoton emitter, we compute the average $N$-photon efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ as
@@
where the probabilities $P_{N}[m]_{y}$ are calculated with the master equation method of Appendix B.1 for each realization $m=1,\dots,M$ of the independent disorder $y$.
The results of these computations for $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ in the case of disorder on the antenna couplings $\Delta\gamma$, as well as in the decoherence rates $\Delta\gamma_{\phi}$ and $\Delta\gamma_{\rm loss}$, are shown in Fig. (8). We perform calculations up to $N=4$ slightly inhomogeneous photon sources, for parameters A and C, and we do not observe any detrimental effect up to a disorder strength of $10\%$ of the average values [cf. Fig. 8(a)-(b)]. This is not surprising in the case of inhomogeneous decoherence rates $\gamma_{\phi}^{j}$ and $\gamma_{\rm loss}^{j}$ since their effects have nothing to do with the synchronization dynamics and are very marginally small anyways $\gamma_{\phi}^{j},\gamma_{\rm loss}^{j}\ll\gamma$. In the case of inhomogeneous antenna couplings $\gamma_{j}=\gamma+\delta\gamma_{j}$, they imply slightly different emission time scales for the emitted single-photons,
@@
Figure 8: Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{y}$, as a function of disorder in antenna decay $y=\Delta\gamma/\gamma$, qubit dephasing $y=\Delta\gamma_{\phi}/\gamma_{\phi}$, and waveguide loss $y=\Delta\gamma_{\rm loss}/\gamma_{\rm loss}$. (a)-(b) Up to a disorder strength of $10\%$, we do not observe any impact of these three disorder sources on the synchronization and demultiplexing dynamics, as calculated up to $N=4$ for parameter sets A and C, respectively. All results are obtained from averaging $M=10^{3}$ realizations for each type of disorder.
On the other hand, inhomogeneities in the qubit frequencies $\delta\omega^{j}_{q}$ and in the couplings $\delta g_{j}$ are much more harmful to the multiphoton synchronization performance and therefore they need to be controlled more precisely in an experimental implementation of the device. Inhomogeneous couplings $g_{j}$, in particular, induce different cavity-mediated driving strengths $\Omega^{j}_{\rm cm}=\Omega_{\rm cm}+|\alpha_{\rm ss}|\delta g_{j}$, and therefore different times to realize an exact $\pi$-pulse on each qubit $\sim\pi/(2\Omega_{\rm cm}^{j})$. Since we control only the global duration of the cavity pulse, we optimally set it to the average $\pi$-pulse time $T=\pi/(2\Omega_{\rm cm})$, but this unavoidably leads to slightly different probabilities $P_{e}^{j}$ of preparing the excited states $|e\rangle_{j}$ on each qubit. Explicitly, we have
@@
and therefore a low coupling disorder $\delta g_{j}\ll g$ is required for a high performance of the synchronized multiphoton emitter.
To precisely quantify the deviation from the ideal photon independence condition, we define the average dependence or demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{y}$ over disorder $y$ as
@@
with ${\cal P}_{1}^{j}[m]_{y}$ calculated from $m=1,\dots,M$ realizations of the disorder $y$ and for each of the $j=1,\dots,N$ SPSs in the setup. In Figs. 9(a)-(b) we compute $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=4$ for parameter set A, and disorder strengths $\Delta g/g=0\%$ (blue), $\Delta g/g={\color[rgb]{0,0,0}1}\%$ (red), and $\Delta g/g={\color[rgb]{0,0,0}2.5}\%$ (green). To achieve reasonable statistical errors $\Delta D_{N}\lesssim 10^{-{\color[rgb]{0,0,0}5}}$, these calculations require $\color[rgb]{0,0,0}M\sim 1.7\times 10^{5}$} realizations. We observe that the $N$-photon efficiency and dependence, $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$, are minimally modified for $\delta g_{j}/g\lesssim{\color[rgb]{0,0,0}2.5\%}$. For the average coupling of $g/2\pi=4$ MHz, considered in parameter sets A and B, suppressing these disorder effects thus requires coupling inhomogeneities in the range $\delta g_{j}/2\pi\sim 0.1$ MHz. This level of inhomogeneity has already been reported in superconducting circuits Wang et al. (2020) and may be further reduced by using symmetric ring cavity configurations Huang et al. (2021); Hazra et al. (2021). For parameter sets C and D, however, reaching the same level of precision requires a coupling inhomogeneity down to $\delta g_{j}/2\pi\sim 5$ kHz which has not been achieved with current superconducting technology, but may be possible in the future by reducing residual stray capacitances in the circuit and also considering symmetric cavity configurations.
Figure 9: Impact of inhomogeneities in qubit-cavity couplings $\delta g_{j}$ and qubit frequencies $\delta\omega_{j}$ on the performance of the synchronized multiphoton emitter. (a) Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ up to $N=4$ for coupling disorder strengths $\Delta g/g=0$ (blue), $\Delta g/g={\color[rgb]{0,0,0}1}\%$ (red), and $\Delta g/g={\color[rgb]{0,0,0}2.5}\%$ (green). (b) Average demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=3$, for the same disorder values as in (a). (c)-(d) Analogous calculations as in (a)-(b), but considering $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ with disorder in the qubit frequencies $\delta\omega^{j}_{q}$. We consider disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green). All calculations are performed using the parameter set A of Table 1 and up to $M=1.7\times 10^{5}$ realizations of each type of disorder.